The following graph shows Bob’s salary from the year 2000 to the year 2003. He was hired in the year 2000; therefore, x = 0 represents the year 2000.
2000=30,000
2001=32,000
2002=34,000
2003=36,000

a) List the coordinates of two points on the graph in (x, y) form.
(___, __),(___, ___)
b) Find the slope of this line:

c) Find the equation of this line in slope-intercept form.

d) If Bob’s salary trend continued, what would his salary be in the year 2005?

Answer by feliz1965(151)

The following graph shows Bob’s salary from the year 2000 to the year 2003. He was hired in the year 2000; therefore, x = 0 represents the year 2000.
2000=30,000
2001=32,000
2002=34,000
2003=36,000
a) List the coordinates of two points on the graph in (x, y) form.
(0, 2000) (3, 2003)
b) Find the slope of this line: The slope is 1.
c) Find the equation of this line in slope-intercept form.
The equation is y = x + 1997
d) If Bob’s salary trend continued, what would his salary be in the year 2005?
We can use the linear function y = 2000x + 30,000 to find Bob's salary in 2005.
From 2005 to 2000, there are 5 years. So, let x = 5.
y = 2000(5) + 30,000
y = 10,000 + 30, 000
y = $40,000